OpenStudy (anonymous):

Which of the binomials below is a factor of this trinomial? 3x2 + 19x + 6

OpenStudy (cwrw238):

3x2 + 19x + 6 = (3x + 1)(x + 6)

OpenStudy (anonymous):

How did you get that?

OpenStudy (cwrw238):

to get 3x^2 you have to have (3x )(x ) now for the other numbers in the parentheses we need 2 numbers whose product is + 6 and when one of them is multiplied by 3 and added to the other gives + 19 these are +6 and + 1 1 + 6 = 6 and 3*6 + 1 = 19

OpenStudy (cwrw238):